OpenStudy (anonymous):

Which of the following points lies on the circle whose center is at the origin and whose radius is 10? (-6, 4) (, ) (6, -8)

OpenStudy (anonymous):

the second one is \[\sqrt{10},\sqrt{10}\]

OpenStudy (lgbasallote):

x^2 + y^2 = 100 plug in each coordinate and see which sums up to 100

OpenStudy (anonymous):

here is the standard form of a circel with center (0,0) and radius 10: x^2+y^2=100, test out your points , see which one give you 100=100

OpenStudy (lgbasallote):

basically trial and error

OpenStudy (anonymous):

basically

OpenStudy (anonymous):

ok ill try

OpenStudy (anonymous):

would it be sqrt of 10 then